# Few-shot Tuning of Foundation Models for Class-incremental Learning

### v1

@@ 4.3 Results
To further evaluate the efficacy of CoACT, we investigate the performance with a different number of samples per class, including 1, 2, 4, 8, and 16. As seen in Figure 8, CoACT outperforms other methods in all settings and shows robustness in very-low-shot settings. The baselines vary in effectiveness with the different number of samples per class. For instance, prototype learning performs better than linear tuning and LoRA in 1-shot and 2-shot settings, while the other two perform better as the number of samples increases. These results indicate that linear tuning and LoRA do not learn very effective representations in very-low-shot settings, compared to our method and prototype learning.
Finally, in Table 2, we present per-dataset results using three different encoder backbones. With ViT-B/16 as the backbone, while CoACT shows average improvements of 2.66%, 2.54%, and 2.47% improvement over prototype learning, linear tuning, and LoRA, respectively. Notably, it shows up to a 12.79% improvement on some individual datasets such as Resisc-45. In general, we find relatively higher improvements for more challenging datasets. For instance, on the 5 datasets with the least accuracy (Country211, FGVCAircraft, GTSRB, Resisc-45, and StanfordCars), CoACT shows an average improvement of 6.14% over the best baseline method (LoRA). We observe similar improvements in the performance with the other two encoders (ViT-B/32 and ViT-L/16), where CoACT outperforms the best baseline (LoRA) by average accuracies of 2.08% and 1.67%.
To further evaluate the effectiveness of our method, we also investigate its performance in the traditional training setup of FSCIL. Following the existing literature, we present these results on the CIFAR-100, CUB-200, and miniImageNet datasets. In Table 3, we compare the performance of CoACT with prior works on the CIAFR-100 dataset, where we divide the existing methods into two groups: the first group, which trains randomly initialized models and the second group which use a pre-trained encoder. An extension of this table with more prior works is presented in the Appendix A.3. As we observe in the table, CoACT outperforms the previous SOTA without a pre-trained encoder (BOT [39]) by 25.88% and the SOTA with a pre-trained encoder (CPE-CLIP [9]) by 4.11%. For a fair comparison with existing methods, we evaluate SAVC [42] and BOT [39] with the same pre-trained encoder (ViT-B/16) as ours, where both methods show considerably lower performance compared to CoACT. Results on the CUB-200 and miniImageNet datasets are presented in Figure 8 (See Appendix A.3). For the CUB-200 dataset, the proposed method outperforms BOT and CPE-CLIP by 17.44% and 5.02%, respectively, while the improvements over these two SOTA methods on the miniImageNet dataset are 36.67% and 13.47%.
Table 3: Comparison to prior works on traditional FSCIL on CIFAR-100 with 60 base classes and 5-way 5-shot incremental setting. We reproduce existing SOTA methods (marked with *) with the same backbone as CoACT. Results on the CUB-200 and miniImageNet are presented in the Appendix.
Method Pre-trained Enc. Acc. in each session (%) ↑
0 1 2 3 4 5 6 7 8
CEC [48] 73.07 68.88 65.26 61.19 58.09 55.57 53.22 51.34 49.14
MetaFSCIL [7] 74.50 70.10 66.84 62.77 59.48 56.52 54.36 52.56 49.97
CLOM [53] 74.2 69.83 66.17 62.39 59.26 56.48 54.36 52.16 50.25
C-FSCIL [15] 77.47 72.40 67.47 63.25 59.84 56.95 54.42 52.47 50.47
LIMIT [51] 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23
FACT [50] 74.60 72.09 67.56 63.52 61.38 58.36 56.28 54.24 52.10
SAVC [42] 79.85 73.70 69.37 65.28 61.91 59.27 57.24 54.97 53.12
SoftNet [21] 79.88 75.54 71.64 67.47 64.45 61.09 59.07 57.29 55.33
BOT [39] 80.25 77.20 75.09 70.82 67.83 64.86 62.73 60.52 58.75
SAVC* [42] ViT-B/16 82.98 75.35 74.01 71.51 70.64 69.78 68.98 67.84 66.24
BOT* [39] ViT-B/16 83.75 78.14 76.85 73.23 72.95 72.03 71.56 70.66 69.72
SV-T [35] SwinT 86.77 82.82 80.36 77.20 76.06 74.00 72.92 71.68 69.75
CPE-CLIP [9] CLIP-B/16 87.83 85.86 84.93 82.85 82.64 82.42 82.27 81.44 80.52
CoACT (Ours) ViT-B/16 90.46 88.46 88.11 86.94 86.98 86.52 86.39 86.0 84.63
Table 4: Performance on taditional FSCIL with a few-shot base session.
Method Base shots Acc. in each session (%) ↑
0 1 2 3 4 5 6 7 8
BOT [39] 500 80.25 77.20 75.09 70.82 67.83 64.86 62.73 60.52 58.75
CPE-CLIP[9] 500 87.83 85.86 84.93 82.85 82.64 82.42 82.27 81.44 80.52
CoACT (Ours) 1 59.04 57.79 58.91 58.65 59.46 59.64 60.01 60.42 59.91
4 72.9 71.11 71.39 70.75 71.24 70.95 71.2 71.16 70.57
8 76.53 74.78 74.71 73.95 74.34 73.89 73.99 73.81 73.05
32 83.87 82.0 81.61 80.24 80.27 79.64 79.56 79.25 78.35
128 83.93 83.1 82.69 82.21 81.94 81.54 81.16 80.90 80.69
Next, we investigate how CoACT compares to previous SOTA methods, namely BOT and CPE-CLIP, in terms of forgetting the already learned classes. As we find from Figure 8, CoACT shows a lower forgetting than previous SOTA methods over the incremental sessions. Specifically, CoACT shows a forgetting of 5.8% compared to the 7.3% of CPE-CLIP and 21.5% of BOT. This further shows the robustness of CoACT against forgetting during incremental learning sessions.
ACT C.F. C.I.T Acc
61.43
61.31
61.28
60.26
58.77
Table 5: Ablation study.
We also study how CoACT performs with a reduced number of samples per base class in the traditional FSCIL setup. For this experiment, we keep the incremental session setup the same (5-way 5-shot for CIFAR-100) but use only a few samples per class for the base session. The results of this study are presented in Table 4. As we find from this experiment, with only 1 sample per base class, CoACT achieves an accuracy of 59.91% in the final session, outperforming BOT with 500 samples per class. We also outperform CPE-CLIP with an accuracy of 80.69% in the final session when we use 128 samples per base class.
@@ 4.4 Ablation study
We present an ablation study on the proposed components of CoACT in Table 5, where ACT, C.F., and C.I.T stand for asynchronous contrastive tuning, controlled fine-tuning, and consistency-guided incremental tuning, respectively. Given that the asynchronous contrastive tuning component of our method could not be removed as it contains the trainable parameters, we start this study by removing controlled fine-tuning and consistency-guided incremental tuning modules individually and simultaneously. Interestingly, we observe that while individual removal of these components does not show considerable drops in performance, their concurrent application within our framework results in a significant boost in performance of 1.17% across 16 datasets. Finally, with the ablation of all three components and only training a linear classifier, we observe a 2.66% drop in performance.
@@
Next, we present a comprehensive study on different parameters of CoACT in Table 6. In the first study (Table 6(a)), we study different alternates for LoRA, namely Adapter [18] and Prompt [27], where we observe that Adapter and Prompt do not perform as well as LoRA, with a final accuracy of 60.01% and 59.87% versus 60.15%, respectively. In the next table (Table 6(b)), we study the performance for different numbers of training epochs, where we find the best results when trained for 50 epochs. Next, in Table 6(c), we study the performance when adding different numbers of LoRA layers to the pre-trained model, where we observe that 12 blocks achieve slightly better results. In the next study (Table 6(d)), we investigate the performance of our method by adding learnable LoRA layers to both encoders, effectively incorporating synchronous training. This variant of our method shows a drop of 1.37% in the final accuracy, showing the importance of our proposed asynchronous teacher. We then study different parameters for the controlled tuning step, specifically the learning rate factor $C_{f}$, fine-tuning layers $C_{l}$, and fine-tuning start epoch $E_{c}$, in Table 6(e-g). The results from this study show the best performance when we fine-tune the pre-trained layers with a higher LR factor, effectively fine-tuning the pre-trained layers with lower LR and allowing smaller changes in the pre-trained weights. For $C_{l}$, we find the best results when we fine-tune half of the layers (6 out of 12) of the pre-trained encoder; training just the final layer or the whole network results in reduced accuracy. Moreover, the best result is obtained when starting the fine-tuning step after 10 epochs of training of the newly added LoRA layers. Lastly, In Table 6(h), we show the results for training different layers during incremental learning where we find the best results for tuning both the classifier and the LoRA layers during incremental tuning.
@@
Classifier (clf) 61.22
clf + LoRA 61.28
LoRA 60.94
Table 6: Ablation of different parameters of CoACT on FSCIL averaged across 16 datasets. Full experiments on all datasets are presented in the Appendix.
@@ 5 Conclusion
To enable few-shot class-incremental learning with pre-trained large vision models, we propose CoACT. Our method can effectively tune a foundation model to learn new classes without losing the generalization of the pre-training or forgetting previously learned classes. Extensive studies show the effectiveness of our method, achieving higher accuracy, lower forgetting, and robustness in low-shot settings. CoACT also outperforms prior SOTA in the standard FSCIL setup by a large margin. We present comprehensive experiments on different components of CoACT and make our code available to foster rapid developments in the area.

### v2

@@ 4.3 FSCIT results
To further evaluate the efficacy of CoACT, we investigate the performance with a different number of samples per class. Specifically, we explore fine-tuning the foundation model in the FSCIT setup with only 1, 2, 4, 8, and 16 samples per class. The results of this experiment are presented in Figure 5. As we find from this experiment, CoACT outperforms other methods in all settings and shows robustness in very low-shot settings. The baselines vary in effectiveness with the different number of samples per class. For instance, prototype learning performs better than linear tuning and LoRA in 1-shot and 2-shot settings, while the other two perform better as the number of samples increases. These results indicate that linear tuning and LoRA do not learn very effective representations in very low-shot settings compared to our method and prototype learning.
Table 7: Ablation study on all datasets. Here, Asy. Cont., Cal. Tun., Con. tun. refers to asynchronous contrastive tuning, controlled fine-tuning and consistency-guided incremental tuning. The reported values are average and standard deviations across three runs.
Asy. Cont.
Cal. Tun.
Con. Tun.
Caltech101
CIFAR100
Country211
CUB200
DTD
EuroSat
FGVCAircraft
Food101
$86.86{\scriptstyle\pm 0.03}$ $78.31{\scriptstyle\pm 0.02}$ $7.42{\scriptstyle\pm 0.04}$ $77.38{\scriptstyle\pm 0.01}$ $55.11{\scriptstyle\pm 0.05}$ $62.25{\scriptstyle\pm 0.03}$ $18.98{\scriptstyle\pm 0.02}$ $71.59{\scriptstyle\pm 0.08}$
$85.89{\scriptstyle\pm 0.05}$ $78.45{\scriptstyle\pm 0.03}$ $7.45{\scriptstyle\pm 0.02}$ $77.27{\scriptstyle\pm 0.04}$ $54.59{\scriptstyle\pm 0.01}$ $62.91{\scriptstyle\pm 0.05}$ $18.20{\scriptstyle\pm 0.03}$ $71.83{\scriptstyle\pm 0.07}$
$86.01{\scriptstyle\pm 0.02}$ $78.96{\scriptstyle\pm 0.01}$ $7.46{\scriptstyle\pm 0.03}$ $77.20{\scriptstyle\pm 0.05}$ $54.82{\scriptstyle\pm 0.02}$ $62.31{\scriptstyle\pm 0.01}$ $18.91{\scriptstyle\pm 0.09}$ $71.47{\scriptstyle\pm 0.03}$
$84.62{\scriptstyle\pm 0.12}$ $78.37{\scriptstyle\pm 0.04}$ $7.49{\scriptstyle\pm 0.01}$ $76.44{\scriptstyle\pm 0.03}$ $54.83{\scriptstyle\pm 0.05}$ $62.22{\scriptstyle\pm 0.02}$ $18.23{\scriptstyle\pm 0.01}$ $71.39{\scriptstyle\pm 0.04}$
$83.47{\scriptstyle\pm 0.10}$ $75.58{\scriptstyle\pm 0.02}$ $7.42{\scriptstyle\pm 0.05}$ $74.10{\scriptstyle\pm 0.01}$ $54.72{\scriptstyle\pm 0.03}$ $62.20{\scriptstyle\pm 0.04}$ $15.51{\scriptstyle\pm 0.02}$ $71.09{\scriptstyle\pm 0.05}$
Asy. Cont.
Cal. Tun.
Con. Tun.
MiniImageNet
Flowers102
OxfordPets
Resisc-45
StanfordCars
SUN397
VOC 2007
GTSRB
Average
$95.34{\scriptstyle\pm 0.03}$ $98.18{\scriptstyle\pm 0.02}$ $87.79{\scriptstyle\pm 0.01}$ $62.62{\scriptstyle\pm 0.04}$ $24.40{\scriptstyle\pm 0.05}$ $66.11{\scriptstyle\pm 0.03}$ $64.45{\scriptstyle\pm 0.02}$ $26.05{\scriptstyle\pm 0.03}$ 61.43
$95.04{\scriptstyle\pm 0.01}$ $98.01{\scriptstyle\pm 0.03}$ $88.25{\scriptstyle\pm 0.02}$ $61.77{\scriptstyle\pm 0.05}$ $24.14{\scriptstyle\pm 0.01}$ $66.01{\scriptstyle\pm 0.04}$ $64.89{\scriptstyle\pm 0.03}$ $26.29{\scriptstyle\pm 0.04}$ 61.31
$95.34{\scriptstyle\pm 0.02}$ $98.15{\scriptstyle\pm 0.01}$ $87.92{\scriptstyle\pm 0.03}$ $61.31{\scriptstyle\pm 0.02}$ $24.07{\scriptstyle\pm 0.04}$ $65.98{\scriptstyle\pm 0.05}$ $64.59{\scriptstyle\pm 0.01}$ $26.05{\scriptstyle\pm 0.02}$ 61.28
$95.34{\scriptstyle\pm 0.04}$ $98.09{\scriptstyle\pm 0.05}$ $88.10{\scriptstyle\pm 0.01}$ $57.21{\scriptstyle\pm 0.03}$ $23.18{\scriptstyle\pm 0.02}$ $65.94{\scriptstyle\pm 0.01}$ $64.72{\scriptstyle\pm 0.04}$ $18.02{\scriptstyle\pm 0.09}$ 60.26
$95.31{\scriptstyle\pm 0.05}$ $98.09{\scriptstyle\pm 0.02}$ $88.10{\scriptstyle\pm 0.04}$ $49.83{\scriptstyle\pm 0.01}$ $18.90{\scriptstyle\pm 0.03}$ $65.68{\scriptstyle\pm 0.02}$ $64.72{\scriptstyle\pm 0.05}$ $15.53{\scriptstyle\pm 0.08}$ 58.77
@@ 4.4 Ablation study
We present an ablation study on the proposed components of CoACT in Table 7. Given that the asynchronous contrastive tuning component of our method could not be removed as it contains the trainable parameters, we start this study by removing controlled fine-tuning and consistency-guided incremental tuning modules individually and simultaneously. Interestingly, we observe that while individual removal of these components does not show considerable drops in performance, their concurrent application within our framework results in a significant boost in performance of 1.17% across 16 datasets. Finally, with the ablation of all three components and only training a linear classifier, we observe a 2.66% drop in performance. Notably, all the ablations experiments show a very small standard deviation.
@@
Next, we present a comprehensive study on different parameters of CoACT in Table 8. In the first study (Table 8(a)), we study different alternates for LoRA, namely Adapter (Houlsby et al., 2019) and Prompt (Jia et al., 2022), where we observe that Adapter and Prompt do not perform as well as LoRA, with a final accuracy of 60.01% and 59.87% versus 60.15%, respectively. In this table, the prompt implementation follows the visual prompt tuning (VPT) (Jia et al., 2022) approach, where learnable prompts are added to the input embeddings of the vision transformer. In the next table (Table 8(b)), we study the performance for different numbers of training epochs, where we find the best results when trained for 50 epochs. Next, in Table 8(c), we study the performance when adding different numbers of LoRA layers to the pre-trained model, where we observe that 12 blocks achieve slightly better results. In the next study (Table 8(d)), we investigate the performance of our method by adding learnable LoRA layers to both encoders, effectively incorporating synchronous training. This variant of our method shows a drop of 1.37% in the final accuracy, showing the importance of our proposed asynchronous teacher. We then study different parameters for the controlled tuning step, specifically the learning rate factor $C_{f}$, fine-tuning layers $C_{l}$, and fine-tuning start epoch $E_{c}$, in Table 8(e-g). The results from this study show the best performance when we fine-tune the pre-trained layers with a higher LR factor, effectively fine-tuning the pre-trained layers with lower LR and allowing smaller changes in the pre-trained weights. For $C_{l}$, we find the best results when we fine-tune half of the layers (6 out of 12) of the pre-trained encoder; training just the final layer or the whole network results in reduced accuracy. Moreover, the best result is obtained when starting the fine-tuning step after 10 epochs of training of the newly added LoRA layers. Table 8(h) shows the results for training different layers during incremental learning, where tuning both the encoder (last few layers) and the LoRA shows the best results. Table 8(i) shows the study on using the encoder from the first session vs. the encoder from the most recent session to enforce consistency. As we find from this table, the student encoder from the first session results in a higher performance. We also study the performance of two new variants of LoRA, including qLoRA (Dettmers et al., 2024) and LoRA+ (Hayou et al., 2024) in Table 8(j). As we find from this table, LoRA+ shows a minor improvement over the Naive LoRA, while qLoRA results in a slight drop in performance. However, in our work, we used the default LoRA method given its popularity and wide use in the area as the default approach to fine-tuning.
@@
Encoder only 61.22
LoRA only 60.94
Enc. + LoRA 61.28
Setup Accuracy
Consistency w/ first 61.43
Consistency w/ last 61.22
Setup Accuracy
LoRA 58.96
qLoRA 58.92
LoRA+ 59.11
Method Param. Thr. (samp./sec.)
Train Infer.
PriViLege 149M 617 1236
CoACT 86M 930 1850
Table 8: Ablation of different parameters of CoACT on FSCIL averaged across 16 datasets, including different alternate for LoRA, training epochs, number of LoRA layers, settings for adding LoRA to the encoders, different parameters for controlled tuning (learning rate factor, fine-tuning layers, and fine-tuning epochs), and different tuning layers.
4.5 Computational complexity
Finally, we analyze the computational complexity of our method compared to the previous SOTA, PriViLege (Park et al., 2024). Our analysis is presented in Table 8(k). In terms of the number of parameters, CoACT has significantly fewer parameters than PriViLege (86 million vs. 149 million) since CoACT is based on a vision-only backbone, whereas PriViLege utilizes a vision-language backbone. Additionally, CoACT achieves higher throughput than PriViLege during both training and inference. Specifically, PriViLege achieves a training throughput of 617 samples/second, whereas CoACT reaches 930. Similarly, the inference throughput is 1,236 for PriViLege and 1,850 for CoACT.
@@ 5 Conclusion
To enable few-shot class-incremental learning with pre-trained large vision models, we propose CoACT. Our method can effectively tune a foundation model to learn new classes without losing the generalization of the pre-training or forgetting previously learned classes. Extensive studies show the effectiveness of our method, achieving higher accuracy, lower forgetting, and robustness in low-shot settings. Notably, CoACT outperforms prior works by up to 5.02% in standard FSCIL setup and by up to 12.51% on FSCIT in the individual datasets. We present comprehensive experiments on different components of CoACT.